3. Suppose G enericFish Corp produced 5,000 cans of herring in the fourth quarter of 2018. GenericFish has two production costs: labor and intermediate goods. GenericFishp its workers S25 per hour, and they employed 1,000 work hours. Intermediate goods were bought from FishyStuff Inc for $4 per can. FishyStuffs only input was labor, which cost $10, 000 total. In the fourth quarter, GenericFish sold 4,000 cans to consumers at a price of $15 ays What was GenericFishs total contribution to GDP in Q4 of 2018? What was FishyStuffs total contribution to national income (wages and profits) in Q4 of 2018? (Round to the nearest whole dollar) A. GDP24:2018 $60,000; NİFishyStuff--$10,000 B. GDP24:2018 : $60,000; NİFishyStuff $20,000 C) GDP24:2018 $69,000; NİFishyStuff $10,000 D. GDPQ4:2018 $69,000; N I FishyStuff $20,000 Consider again the scenario posed in question 3. Suppose in the first quarter of 2019, Gener- cFish produced 5,800 cans. GenericFishs wage per hour went up by 35%. GenericFish employed the same number of work hours. Also, FishyStuffs labor costs were the same as in the previous quarter. GenericFish sold to consumers all cans produced in the first quarter of 2019 and the cans left in inventory from the previous quarter (fourth quarter 2018). However, GenericFish raised the price of their canned fish to $17 after a fishy mar- keting campaign that implied anti-aging benefits associated with the consumption of their product. What were the total wages and profits for Q1 2019? A. ProfitTotul = 62,850; WagesTotul = 43,750 ProfitTotal = 62.850; Wages-Total = 30,000 C. ProfitTotal=54,850; WagesTotal=43,750 Profitrotal 54,850; WagesTotal 39,000
Answer #1

Q.3.

By the output method of GDP calculation,

GenericFish total contribution to GDP = No. of units sold * Price of output

= 4000 * 15 = 60000

National income produced due to Fishy stuff = profits made (by selling intermediate goods for 5000 cans to Generic Fish)

= 5000 * 4 - 10000 (wages paid)

= 20000 - 10000 = 10000

Correct Ans - A